# Scaling limits of critical FK-decorated random planar maps with $q=4$

### v1

@@ Abstract
We establish the first scaling limit for FK($q$)-weighted planar maps in the critical case $q=4$, resolving a problem that has remained open since Sheffield’s seminal work [SheffieldScott2016QGAI]. In that work, Sheffield proved a scaling limit for $q<4$ via the celebrated hamburger-cheeseburger bijection, which initiated the peanosphere (mating-of-trees) approach to Liouville quantum gravity. We prove that, at criticality, the associated burger count $\mathcal{S}$ and discrepancy $\mathcal{D}$ satisfy
@@
where $B^{1}$ and $B^{2}$ are independent two-sided Brownian motions. To the best of our knowledge, no conjecture for the correct discrepancy scaling factor had previously been formulated. Matching the limiting process with the critical mating of trees [AHPS], we establish the first rigorous planar map convergence towards CLE4 and critical ($\gamma=2$) Liouville quantum gravity, in the peanosphere sense. Our proof is based on a novel approach that reveals the exactly solvable nature of the model through a correspondence with the (bicoloured) fully packed loop-$O(2)$ model on triangulations, and yields critical geometric exponents matching the predictions of conformal field theory.
@@ 1 Introduction
Liouville quantum gravity (LQG) surfaces are canonical models of random planar geometry, originally motivated by conformal field theory and string theory [polyakovstrings]. They possess a rich mathematical structure, have deep connections to probability, geometry, combinatorics and theoretical physics [knizhnik1988fractal, DuplantierSheffield, DKRV], and remain the subject of many open problems; see [BerestyckiNathanaël2024Gffa] for an extensive exposition. A central conjecture is that LQG surfaces, which form a one-parameter family indexed by $\gamma\in(0,2]$, arise as scaling limits of discrete random surfaces known as random planar maps.
The best known results in this direction concern models where face degrees are suitably controlled. Progress culminated in the works of Le Gall [le2013uniqueness], Miermont [miermont2013brownian], Miller–Sheffield [LQGTBM] and Holden–Sun [HoldenSun], which establish convergence in the uniform bounded-degree case ($\gamma=\sqrt{8/3}$). It has been predicted in physics, see for example [DKRV] for a precise conjecture, that a natural way to escape from the pure-gravity universality class is to couple the maps with critical statistical mechanics models. Proving rigorous convergence results in this general setting remains a major open problem, although spectacular progress has been made recently for the related model of Boltzmann planar maps with heavy-tailed face degree distributions [curien2025scaling].
@@
where $\text{K}(\mathfrak{p})$, $\text{E}(\mathfrak{p})$ and $\text{V}(\mathfrak{m})$ denote, respectively, the number of connected components, edges and vertices of $\mathfrak{p}$ or $\mathfrak{m}$. Equivalently, the model can be constructed by first sampling a planar map weighted by the self-dual FK($q$)-partition function, and then drawing an FK($q$)-percolation configuration on top of it [duminil2013parafermionic]. The infinite-volume model is obtained as the local limit of such maps as the size tends to infinity; see [SheffieldScott2016QGAI, ChenLinxiao2017Bpot].
This planar map model is intrinsically related to the $q$-state Potts, Ising, and loop-$O(n)$ models. On a fixed lattice, it has been shown in the landmark paper [duminil2021discontinuity] that the model displays a phase transition at $q=4$, while it is known that on planar maps some of the observables degenerate at this critical threshold, see [SheffieldScott2016QGAI, feng2023triviality]. This makes the boundary case $q=4$ particularly challenging and significant, as fundamentally new geometric and probabilistic behaviour emerges.
@@
A powerful approach to studying these maps comes from bijections due to Mullin [mullin67] (for $q=0$), and Bernardi [bernardi08] followed by Sheffield [SheffieldScott2016QGAI] for general $q\geq 0$. These bijections encode (infinite) FK-decorated planar maps by (bi-infinite) words in a finite alphabet, interpreted as an inventory process of two item types (“hamburgers” and “cheeseburgers” in the terminology of [SheffieldScott2016QGAI]). The resulting word defines a (bi-infinite) two-dimensional lattice path, which encodes the geometry of the underlying planar map.
In [SheffieldScott2016QGAI], Sheffield proved that these non-Markovian bi-infinite lattice paths, when rescaled by $n^{-1/2}$, converge to a correlated planar Brownian motion when $q<4$. Via the remarkable mating-of-trees theory [DuplantierMillerSheffield], this proves convergence of FK-decorated maps to $\gamma$-LQG surfaces decorated by $\mathrm{CLE}_{16/\gamma^{2}}$ in the so-called peanosphere topology. This is a powerful result in its own right, and an important step towards resolving the conjecture for conformally embedded maps, see for instance [tutteLQG].
@@
as $n\to\infty$, so a different rescaling is needed to observe non-trivial behaviour in $\mathcal{D}_{n}$. Identifying the correct scaling and obtaining a limit was posed as an open problem in [SheffieldScott2016QGAI]. Our main result fully answers this question.
@@
We also determine the asymptotic variance of $\mathcal{D}_{n}$, which was only shown to be $o(n)$ in [SheffieldScott2016QGAI]. We emphasise that our proof of 1.1 does not rely on this variance estimate.
@@
Our results gain further significance in light of recent work showing that critical LQG surfaces and $\mathrm{CLE}_{4}$ can be represented via mating-of-trees encodings using planar Brownian motion: [AHPS]. Moreover, our techniques allow us to obtain scaling exponents for geometric observables in the FK(4)(4)-planar map model.
For this, we need the notion of a typical loop $\mathfrak{L}$, a typical cluster $\mathfrak{c}$ and a typical envelope $\mathfrak{e}$ in the infinite FK(4)(4)-planar map, as studied for $q<4$ in [berestycki2017critical, GwynneEwain2019Slft] (the envelope is also referred to as a bubble in [berestycki2017critical]). These are defined precisely in Section 2.3 and correspond to choosing a loop, cluster or envelope uniformly at random from an FK(4)(4)-planar map of finite size $k$, and letting $k\to\infty$. We give natural definitions of the length $|\mathfrak{L}|$ of a typical loop, the boundary length $|\partial\mathfrak{c}|$ of a typical cluster and the boundary length $|\partial\mathfrak{e}|$ of a typical envelope in Section 2.3.
@@
The exponents are consistent with the results of [berestycki2017critical, GwynneEwain2019Slft], taking a limit as $q\to 4$, although no conjecture was made for the explicit slowly varying corrections. We point out that in contrast to the previous literature [berestycki2017critical, GwynneEwain2019Slft], our strategy involves exact calculations (see below) that allow us to identify these exponents, which we do before establishing the scaling limit of Theorem 1.1. In the subcritical case, the scaling limit of $(\mathcal{C}_{n},\mathcal{H}_{n})$ is the main tool used to obtain the exponents. We also stress that, contrary to [berestycki2017critical, GwynneEwain2019Slft], our asymptotics are exact and do not involve any $\ell^{o(1)}$ correction.
We remark that it is the observable $|\partial\mathfrak{e}|$ that can be identified most easily with a continuum counterpart. Namely, in the works [lqgcle4], it is shown that exploring a critical Liouville quantum gravity surface and independent conformal loop ensemble with parameter $4$ in an appropriate way, the quantum lengths of the components separated from a quantum typical point are precisely given by the jumps of a $1$-stable process, i.e. a process whose intensity scales the same way as the density of $|\partial\mathfrak{e}|$. We give a more precise interpretation in Section 5.2, which requires a little more background.
The first key step in our strategy is to use a correspondence between the FK(4)(4) decorated map model and the fully packed loop-$O(2)$ model on triangulations, see Section 2.4. For the latter model, using powerful tools from analytic combinatorics developed by [borot2012recursive], we obtain an exact expression and asymptotic identity for the partition function, which is a novel result of independent interest, making rigorous the physics prediction [gaudin1989n].
@@
Note that the form of the asymptotic behaviour features the exponent $2$, which is indicative of criticality. Previously, the asymptotics of the $O(2)$ partition function were obtained only for the “rigid” model [aidekon2024scaling, kammerer2024gaskets], using a different probabilistic argument. We also mention that it is partly possible to extend the results to FK(q)(q) planar maps with $q\in(0,4)$. In a concurrent joint work with Berestycki, the first-named author pursues this approach to derive exact tail exponents (analogous to 1.3) and an explicit expression and asymptotic for the partition function of the associated $O(n)$ model (as in 1.3). However, we note that this analysis builds on Sheffield’s convergence result [SheffieldScott2016QGAI].
Our strategy to establish the scaling limit result (1.1) proceeds as follows. We begin by mapping the FK(4)(4)-weighted planar map model to the fully-packed $O(2)$ model on triangulations. Within this framework, we derive an exact expression for the partition function $F_{\ell}$ using tools from analytic combinatorics (1.4). We then reveal a natural probabilistic interpretation of this partition function in terms of two random-walk-type observables that encode complementary peeling explorations on the map, which we refer to as the exploration into the past and the exploration into the future. We rigorously establish the scaling limits of these two explorations, and, in the course of this analysis, derive 1.3 from the study of the former. Finally, we use ideas from excursion theory to “squeeze” the discrepancy $\mathcal{D}_{n}$ between the two explorations and thereby deduce the full scaling limit from their individual scaling limits. We refer to Section 2.5 for a more detailed overview of the proof outline.
@@ Acknowledgements.
We are grateful to Xin Sun for suggesting the article [borot2012recursive] as a starting point for deriving exact partition functions. We thank Élie Aïdékon, Nathanaël Berestycki, Yuyang Feng, Ewain Gwynne and Julie Tourniaire for insightful discussions. W.D.S. acknowledges the support of the Austrian Science Fund (FWF) grant on “Emergent branching structures in random geometry” (DOI: 10.55776/ESP534). X.H. is supported by the China Scholarship Council. The research of E.P. is supported by UKRI Future Leader’s Fellowship MR/W008513. M.D.W. is supported by a start-up allowance from Croucher Foundation and a start-up fund from Faculty of Science, The University of Hong Kong.
@@ 2.1.1 Definition of the model
Notice that the weight on the right hand side does not depend on the choice of root edge in $M_{k}$ and thus, conditionally on the map, the oriented root edge is chosen uniformly at random. Moreover, conditionally on the map $M_{k}$, the edge configuration $P_{k}$ is sampled as a self-dual FK(q)(q)-percolation (also called random cluster) configuration on $M_{k}$, see e.g. [duminil2013parafermionic].
@@ 2.1.2 The Mullin–Bernardi–Sheffield bijection
In this section we recall the bijective frameworks of Mullin [mullin67], Bernardi [bernardi08] and Sheffield [SheffieldScott2016QGAI], which relate random planar maps decorated by edge sets to words in a certain alphabet $\Theta$ (see below), which can be viewed as inventory accumulation processes.
@@
To summarise, we have explained how, starting with $\mathfrak{m}\in\mathcal{M}_{k}$ and $\mathfrak{p}\subset E(\mathfrak{m})$, we can produce a word $w$ in the alphabet $\Theta$. Before discussing the inverse of this operation, we make some remarks on the word $w$. The length (i.e. number of letters) of $w$ is $2k$ since there are $2k$ triangles. It can be seen [SheffieldScott2016QGAI, Section 4.1] that the matches of burgers and orders correspond to the completion of a quadrangle, in the sense that two symbols corresponding to companion triangles are actually matched under the aforementioned reduction rule. [^2] In particular, we claim that symbols corresponding to fictional triangles are properly matched in $w$. Visually, this means that every burger produced inside $\mathfrak{l}_{1}$ (say) will be consumed inside $\mathfrak{l}_{1}$. For this property to hold, it is important that we wait for the last pair of triangles $(t,t^{\prime})$ before entering $\mathfrak{l}_{1}$ in our construction of the space-filling loop $\mathfrak{l}$. As a consequence, the reduced word $\overline{w}$ of $w$ is equal to $\emptyset$. By construction, the number of loops $L(\mathfrak{m},\mathfrak{p})$ is the number of $\mathsf{F}$ letters in $w$ plus 1.
@@ 2.1.3 Random inventory accumulation
Through the bijection discussed above, a random decorated FK planar map with $k$ edges, i.e., a random pair ${M_{k}}\in\mathcal{M}_{k}$ and ${P_{k}}\subset E{M_{k}})$ corresponds to a random word $W$ in $\Theta$ of length $2k$ satisfying $\overline{W}=\emptyset$. The law of this random word can be described as follows.
@@
Now consider a random word $W$ with $2k$ letters, each sampled independently with distribution given by (2.3), but conditioned on $\overline{W}=\emptyset$. Then the probability of choosing a given word $w$ with $\overline{w}=\varnothing$ satisfies
@@ 2.2.1 Definition of the model
In the preceding section, we have seen that a (random) FK planar map $(M_{k},P_{k})$ with $k$ edges and parameter $q$ corresponds to a word $W$ sampled according to (2.4). When $k\to\infty$, it is then proved in [ChenLinxiao2017Bpot] that $(M_{k},P_{k})$ converges in distribution in the local topology to an object $(M_{\infty},P_{\infty})$, called the infinite FK(q)(q) planar map. More precisely, the convergence takes place on the completion of the space of finite rooted planar maps with a distinguished subgraph, equipped with the local metric $d_{\mathrm{loc}}((\mathfrak{m},\mathfrak{p}),(\mathfrak{m}^{\prime},\mathfrak{p}^{\prime}))={1/\sup\{R}:B_{R}((\mathfrak{m},\mathfrak{p}))=B_{R}((\mathfrak{m}^{\prime},\mathfrak{p}^{\prime}))\}$ where $B_{R}(\mathfrak{m},\mathfrak{p})$ is everything in $(\mathfrak{m},\mathfrak{p})$ at graph distance less than $R$ from the root edge.
@@
In such a bi-infinite word $W$, it is the case that with probability one, there is a match for every symbol (see [SheffieldScott2016QGAI, Proposition 2.2]). That is, any letter $X(i)$ in $W$ corresponding to a burger $(\mathsf{c}$ or $\mathsf{h}$) will be consumed by some order ($\mathsf{C},\mathsf{H},\mathsf{F})$ in the future (i.e., after applying the reduction rules to $W$) and any order will be fulfilled by some burger in the past. The unique match of $X(i)$ is denoted by $X(\varphi(i))$.

### v2

@@ Abstract
We establish the first scaling limit for FK($q$)-weighted planar maps in the critical case $q=4$, resolving a problem that has remained open since Sheffield’s seminal work [SHE16]. In that work, Sheffield proved a scaling limit for $q<4$ via the celebrated hamburger-cheeseburger bijection, which initiated the peanosphere (mating-of-trees) approach to Liouville quantum gravity. We prove that, at criticality, the associated burger count $\mathcal{S}$ and discrepancy $\mathcal{D}$ satisfy
@@
where $B^{1}$ and $B^{2}$ are independent two-sided Brownian motions. To the best of our knowledge, no conjecture for the correct discrepancy scaling factor had previously been formulated. Matching the limiting process with the critical mating of trees [AHP+23], we establish the first rigorous planar map convergence towards CLE4 and critical ($\gamma=2$) Liouville quantum gravity, in the peanosphere sense. Our proof is based on a novel approach that reveals the exactly solvable nature of the model through a correspondence with the (bicoloured) fully packed loop-$O(2)$ model on triangulations, and yields critical geometric exponents matching the predictions of conformal field theory.
@@ 1 Introduction
Liouville quantum gravity (LQG) surfaces are canonical models of random planar geometry, originally motivated by conformal field theory and string theory [POL81]. They possess a rich mathematical structure, have deep connections to probability, geometry, combinatorics and theoretical physics [KPZ88, DS11, DKR+16], and remain the subject of many open problems; see [BerestyckiNathanaël2024Gffa] for an extensive exposition. A central conjecture is that LQG surfaces, which form a one-parameter family indexed by $\gamma\in(0,2]$, arise as scaling limits of discrete random surfaces known as random planar maps.
The best known results in this direction concern models where face degrees are suitably controlled. Progress culminated in the works of Le Gall [LE 13], Miermont [MIE13], Miller–Sheffield [MS21] and Holden–Sun [HS23], which establish convergence in the uniform bounded-degree case ($\gamma=\sqrt{8/3}$). It has been predicted in physics, see for example [DKR+16] for a precise conjecture, that a natural way to escape from the pure-gravity universality class is to couple the maps with critical statistical mechanics models. Proving rigorous convergence results in this general setting remains a major open problem, although spectacular progress has been made recently for the related model of Boltzmann planar maps with heavy-tailed face degree distributions [CMR25].
@@
where $\text{K}(\mathfrak{p})$, $\text{E}(\mathfrak{p})$ and $\text{V}(\mathfrak{m})$ denote, respectively, the number of connected components, edges and vertices of $\mathfrak{p}$ or $\mathfrak{m}$. Equivalently, the model can be constructed by first sampling a planar map weighted by the self-dual FK($q$)-partition function, and then drawing an FK($q$)-percolation configuration on top of it [DUM13]. The infinite-volume model is obtained as the local limit of such maps as the size tends to infinity; see [SHE16, CHE17].
This planar map model is intrinsically related to the $q$-state Potts, Ising, and loop-$O(n)$ models. On a fixed lattice, it has been shown in the landmark paper [DGH+21] that the model displays a phase transition at $q=4$, while it is known that on planar maps some of the observables degenerate at this critical threshold, see [SHE16, FEN25]. This makes the boundary case $q=4$ particularly challenging and significant, as fundamentally new geometric and probabilistic behaviour emerges.
@@
A powerful approach to studying these maps comes from bijections due to Mullin [MUL67] (for $q=0$), and Bernardi [BER08] followed by Sheffield [SHE16] for general $q\geq 0$. These bijections encode (infinite) FK-decorated planar maps by (bi-infinite) words in a finite alphabet, interpreted as an inventory process of two item types (“hamburgers” and “cheeseburgers” in the terminology of [SHE16]). The resulting word defines a (bi-infinite) two-dimensional lattice path, which encodes the geometry of the underlying planar map.
In [SHE16], Sheffield proved that these non-Markovian bi-infinite lattice paths, when rescaled by $n^{-1/2}$, converge to a correlated planar Brownian motion when $q<4$. Via the remarkable mating-of-trees theory [DMS21], this proves convergence of FK-decorated maps to $\gamma$-LQG surfaces decorated by $\mathrm{CLE}_{16/\gamma^{2}}$ in the so-called peanosphere topology. This is a powerful result in its own right, and an important step towards resolving the conjecture for conformally embedded maps, see e.g. [GMS21].
@@
as $n\to\infty$, so a different rescaling is needed to observe non-trivial behaviour in $\mathcal{D}_{n}$. Identifying the correct scaling and obtaining a limit was posed as an open problem in [SHE16]. Our main result fully answers this question.
@@
We also determine the asymptotic variance of $\mathcal{D}_{n}$, which was only shown to be $o(n)$ in [SHE16]. We emphasise that our proof of 1.1 does not rely on this variance estimate.
@@
Our results gain further significance in light of recent work showing that critical LQG surfaces and $\mathrm{CLE}_{4}$ can be represented via mating-of-trees encodings using planar Brownian motion: [AHP+23]. Moreover, our techniques allow us to obtain scaling exponents for geometric observables in the FK(4)(4)-planar map model.
For this, we need the notion of a typical loop $\mathfrak{L}$, a typical cluster $\mathfrak{c}$ and a typical envelope $\mathfrak{e}$ in the infinite FK(4)(4)-planar map, as studied for $q<4$ in [BLR17, GMS19] (the envelope is also referred to as a bubble in [BLR17]). These are defined precisely in Section 2.3 and correspond to choosing a loop, cluster or envelope uniformly at random from an FK(4)(4)-planar map of finite size $k$, and letting $k\to\infty$. We give natural definitions of the length $|\mathfrak{L}|$ of a typical loop, the boundary length $|\partial\mathfrak{c}|$ of a typical cluster and the boundary length $|\partial\mathfrak{e}|$ of a typical envelope in Section 2.3.
@@
The exponents are consistent with the results of [BLR17, GMS19], taking a limit as $q\to 4$, although no conjecture was made for the explicit slowly varying corrections. We point out that in contrast to the previous literature [BLR17, GMS19], our strategy involves exact calculations (see below) that allow us to identify these exponents, which we do before establishing the scaling limit of Theorem 1.1. In the subcritical case, the scaling limit of $(\mathcal{C}_{n},\mathcal{H}_{n})$ is the main tool used to obtain the exponents. We also stress that, contrary to [BLR17, GMS19], our asymptotics are exact and do not involve any $\ell^{o(1)}$ correction.
It is the observable $|\partial\mathfrak{e}|$ that can be identified most easily with a continuum counterpart. Namely, in the work [AG25], it is shown that exploring a critical Liouville quantum gravity surface and independent conformal loop ensemble with parameter $4$ in an appropriate way, the components separated from a quantum typical point are conditionally independent critical quantum discs given their boundary lengths. Moreover, it is a consequence of the critical mating of trees [AHP+23] that these boundary lengths can be described as the jumps of a $1$-stable process, i.e. a process whose intensity scales the same way as the density of $|\partial\mathfrak{e}|$. We give a more precise interpretation in Section 5.2, which requires a little more background.
The first key step in our strategy is to use a correspondence between the FK(4)(4) decorated map model and the fully packed loop-$O(2)$ model on triangulations, see Section 2.4. For the latter model, using powerful tools from analytic combinatorics developed by [BBG11], we obtain an exact expression and asymptotic identity for the partition function, which is a novel result of independent interest, making rigorous the physics prediction [GK89].
@@
Note that the form of the asymptotic behaviour features the exponent $2$, which is indicative of criticality. Previously, the asymptotics of the $O(2)$ partition function were obtained only for the “rigid” model [ADH24, KAM24], using a different probabilistic argument. We also mention that it is partly possible to extend the results to FK(q)(q) planar maps with $q\in(0,4)$. In a concurrent joint work with Berestycki [BD25], the first-named author pursues this approach to derive exact tail exponents (analogous to 1.3) and an explicit expression and asymptotic for the partition function of the associated $O(n)$ model (as in 1.3). However, we note that this analysis builds on Sheffield’s convergence result [SHE16].
Our strategy to establish the scaling limit result (1.1) proceeds as follows. We begin by mapping the FK(4)(4)-weighted planar map model to the fully-packed $O(2)$ model on triangulations. Within this framework, we derive an exact expression for the partition function $F_{\ell}$ using tools from analytic combinatorics (1.4). We then reveal a natural probabilistic interpretation of this partition function in terms of two random-walk-type observables that encode complementary peeling explorations on the map, which we refer to as the exploration into the past and the exploration into the future. We rigorously establish the scaling limits of these two explorations, and, in the course of this analysis, derive 1.3 from the study of the former. Finally, we use ideas from excursion theory to “squeeze” the discrepancy $\mathcal{D}_{n}$ between the two explorations and thereby deduce the full scaling limit. We refer to Section 2.5 for a more detailed overview of the proof outline.
@@ Acknowledgements.
We are grateful to Xin Sun for suggesting the article [BBG11] as a starting point for deriving exact partition functions. We thank Élie Aïdékon, Nathanaël Berestycki, Yuyang Feng, Ewain Gwynne and Julie Tourniaire for insightful discussions. W.D.S. acknowledges the support of the Austrian Science Fund (FWF) grant on “Emergent branching structures in random geometry” (DOI: 10.55776/ESP534). X.H. is supported by the China Scholarship Council. The research of E.P. is supported by UKRI Future Leader’s Fellowship MR/W008513. M.D.W. is supported by a start-up allowance from Croucher Foundation and a start-up fund from Faculty of Science, The University of Hong Kong.
@@ 2.1.1 Definition of the model
Notice that the weight on the right hand side does not depend on the choice of root edge in $M_{k}$ and thus, conditionally on the map, the oriented root edge is chosen uniformly at random. Moreover, conditionally on the map $M_{k}$, the edge configuration $P_{k}$ is sampled as a self-dual FK(q)(q)-percolation (also called random cluster) configuration on $M_{k}$, see e.g. [DUM13].
@@ 2.1.2 The Mullin–Bernardi–Sheffield bijection
In this section we recall the bijective frameworks of Mullin [MUL67], Bernardi [BER08] and Sheffield [SHE16], which relate random planar maps decorated by edge sets to words in a certain alphabet $\Theta$ (see below), which can be viewed as inventory accumulation processes.
@@
To summarise, we have explained how, starting with $\mathfrak{m}\in\mathcal{M}_{k}$ and $\mathfrak{p}\subset E(\mathfrak{m})$, we can produce a word $w$ in the alphabet $\Theta$. Before discussing the inverse of this operation, we make some remarks on the word $w$. The length (i.e. number of letters) of $w$ is $2k$ since there are $2k$ triangles. It can be seen [SHE16, Section 4.1] that the matches of burgers and orders correspond to the completion of a quadrangle, in the sense that two symbols corresponding to companion triangles are actually matched under the aforementioned reduction rule. [^2] In particular, we claim that symbols corresponding to fictional triangles are properly matched in $w$. Visually, this means that every burger produced inside $\mathfrak{l}_{1}$ (say) will be consumed inside $\mathfrak{l}_{1}$. For this property to hold, it is important that we wait for the last pair of triangles $(t,t^{\prime})$ before entering $\mathfrak{l}_{1}$ in our construction of the space-filling loop $\mathfrak{l}$. As a consequence, the reduced word $\overline{w}$ of $w$ is equal to $\emptyset$. By construction, the number of loops $L(\mathfrak{m},\mathfrak{p})$ is the number of $\mathsf{F}$ letters in $w$ plus 1.
@@ 2.1.3 Random inventory accumulation
Through the bijection discussed above, a random decorated FK planar map with $k$ edges, i.e., a random pair ${M_{k}}\in\mathcal{M}_{k}$ and ${P_{k}}\subset E(M_{k})$, corresponds to a random word $W$ in $\Theta$ of length $2k$ satisfying $\overline{W}=\emptyset$. The law of this random word can be described as follows.
@@
Now consider a random word $W$ with $2k$ letters, each sampled independently with distribution given by (2.3), but conditioned on $\overline{W}=\emptyset$. Then the probability of choosing a given word $w$ with $\overline{w}=\emptyset$ satisfies
@@ 2.2.1 Definition of the model
In the preceding section, we have seen that a (random) FK planar map $(M_{k},P_{k})$ with $k$ edges and parameter $q$ corresponds to a word $W$ sampled according to (2.4). When $k\to\infty$, it is then proved in [CHE17] that $(M_{k},P_{k})$ converges in distribution in the local topology to an object $(M_{\infty},P_{\infty})$, called the infinite FK(q)(q) planar map. More precisely, the convergence takes place on the completion of the space of finite rooted planar maps with a distinguished subgraph, equipped with the local metric $d_{\mathrm{loc}}((\mathfrak{m},\mathfrak{p}),(\mathfrak{m}^{\prime},\mathfrak{p}^{\prime}))=1/\sup\{R:B_{R}((\mathfrak{m},\mathfrak{p}))=B_{R}((\mathfrak{m}^{\prime},\mathfrak{p}^{\prime}))\}$ where $B_{R}(\mathfrak{m},\mathfrak{p})$ is everything in $(\mathfrak{m},\mathfrak{p})$ at graph distance less than $R$ from the root edge.
@@
In such a bi-infinite word $W$, it is the case that with probability one, there is a match for every symbol (see [SHE16, Proposition 2.2]). That is, any letter $X(i)$ in $W$ corresponding to a burger $(\mathsf{c}$ or $\mathsf{h}$) will be consumed by some order ($\mathsf{C},\mathsf{H},\mathsf{F})$ in the future (i.e., after applying the reduction rules to $W$) and any order will be fulfilled by some burger in the past. The unique match of $X(i)$ is denoted by $X(\varphi(i))$.